OpenStudy (anonymous):

if cosɵ=3x then tan^2ɵ= ?

zepdrix (zepdrix):

\[\Large\rm \cos \theta=\frac{adjacent}{hypotenuse}=\frac{3x}{1}\]You can use this information to draw a triangle: